# Revolutionizing Mixed Precision Quantization: Towards Training-free Automatic Proxy Discovery via Large Language Models

### v1

@@ 3.3 Evolutionary Framework of the TAP
Step 0: Initialization. At time step 0, an (initialization, empty context) prompt is input to the reference LLM ($\mathcal{L}_{\text{ref}}$) to generate the initial proxy population $\mathcal{P}_{0}=\{f_{1},\dots,f_{N}\}$. Each proxy $f_{i}$ must include three components: 1. Natural language reasoning logic for sensitivity scoring (e.g., "convolutional channel sensitivity = L2 norm of weights"); 2. Executable code for computing channel-level/layer-level sensitivity scores; 3. A bit-width allocation function that maps sensitivity scores and target compression rates to specific bit-widths (e.g., 2/4/8 bits for convolutional channels, 4/8 bits for linear layers).
Meanwhile, the LLM to be optimized ($\mathcal{L}_{\theta}$) is initialized with parameters consistent with $\mathcal{L}_{\text{ref}}$.
Step 1: Proxy Generation. At generation $t$, a context window $\mathcal{C}_{t}\subseteq\mathcal{P}_{t}$ (containing the logic and performance of historically high-fitness proxies) is selected from the current population $\mathcal{P}_{t}$. The $\mathcal{C}_{t}$ and operation type prompts (mutation/crossover) are input to $\mathcal{L}_{\theta}$ to generate the candidate proxy set $\mathcal{P}_{t}^{\prime}$—the sensitivity scoring logic and bit-allocation rules of these proxies have been adjusted based on previous DPO optimization results.
@@
Step 3: DPO Strategy Update. 1. Construct preference data pairs: Sort the proxies in $\mathcal{P}_{t}\cup\mathcal{P}_{t}^{\prime}$ by $\phi(f)$ and pair them two-by-two to generate the preference dataset $\mathcal{D}=\{(f_{\text{prefer}},f_{\text{disprefer}})\}$; 2. Optimize LLM parameters: Input $\mathcal{D}$ into the DPO module, and update the parameters of $\mathcal{L}_{\theta}$ by minimizing the core optimization formula in Section 3.2, enabling the prompts generated by the LLM to more easily induce high-fitness proxies; 3. Update the context window: Add the logic and performance of the top-10% high-fitness proxies in $\mathcal{P}_{t}\cup\mathcal{P}_{t}^{\prime}$ to the context library, providing better references for the next generation of proxy generation.
@@
PACT [choi2018pact] 69.8 {5}/{5} 5/5 7.2 - 83.86 Fixed-precision
LSQ [bhalgat2020lsq+] 67.6 {2}/{2} 2/2 - - - Fixed-precision
PDQ [chu2019mixed] 65.0 {1,2,4,8}/{2} -/- - - - Fixed-precision
Hybrid-Net* [chakraborty2020constructing] 62.7 {2}/{4} $\ast$$\ast - - -$ Post-training Deterministic
HAQ [wang2019haq] 70.4 [2,8]/{32} $\ast$/32 5.8 - 87.47 RL-based
HAWQ [dong2019hawq] 68.5 {2}/{4} $\ast$$\ast -$ 15.6 - Sensitivity
HAWQ-V3 [yao2021hawq] 70.4 {4}/{8} $\ast$$\ast$ 7.0 - 84.31 Sensitivity
DNAS [wu2018mixed] 70.0 {1,2,4,8,32}/{1,2,4,8,32} -/- 6.8 - 84.75 Differentiable
EdMIPS [cai2020rethinking] 65.9 [1,4]/[2,4] $\ast$$\ast$ 4.7 9.5 89.46 Differentiable
EdMIPS-C [cai2020rethinking] 59.1 [1,4]/[2,4] $\ast$$\ast$ 4.5 0.6 89.91 Differentiable
FracBits-SAT [yang2021fracbits] 70.6 [2,8]/[2,8] $\ast$$\ast$ 5.81 - 86.97 Differentiable
GMPQ-C [wang2021generalizable] 69.9 [2,8]/[2,8] $\ast$$\ast$ 4.1 0.6 90.81 Differentiable
SMPQ [kang2025and] 72.6 [2,8]/{8} $\ast$/8 - 4.9 - Differentiable
ASGA [ma2025learning] 66.4 [2,8]/[2,8] $\ast$$\ast -$ 0.6 - Differentiable
OMPQ [ma2023ompq] 72.08 [2,8]/{5} $\ast$/5 6.7 0.45 84.98 Training-free
EMQ [dong2023emq] 72.28 [2,8]/{6} $\ast$/6 6.7 0.51 84.98 Training-free
TAP-C 72.93 [2, 8]/{8} $\ast$/8 7.5 $9.17\times 10^{-6}$ 83.18 Training-free
@@
PACT [choi2018pact] 76.70 {5}/{5} 4/4 16.0 - 83.64 Fixed-precision
HAQ [wang2019haq] 75.48 [2,8]/{32} $\ast$/32 9.62 - 90.16 RL-based
LQ-net [zhang2018lq] 76.40 {4}/{32} 4/32 13.1 - 84.7 Differentiable
HAWQ-V3 [yao2021hawq] 75.39 {4}/{8} $\ast$$\ast$ 18.7 - 78.1 Differentiable
@@
Onebit-width [koryakovskiy2023one] 76.70 [2,8]/{8} $\ast$/8 12.3 - 87.42 Differentiable
SMPQ [kang2025and] 76.2 [2,8]/[2,8] $\ast$$\ast$ 12.4 - 87.32 Differentiable
ASGA [ma2025learning] 71.5 [2,8]/[2,8] $\ast$$\ast -$ 0.6 - Differentiable
OMPQ [ma2023ompq] 76.28 [2,8]/{5} $\ast$/5 18.7 0.46 80.88 Training-free
EMQ [dong2023emq] 76.70 [2,8]/{6} $\ast$/6 17.9 0.52 81.70 Training-free
TAP-C 76.81 [2,8]/{8} $\ast$$\ast$ 18.6 $8.75\times 10^{-6}$ 80.98 Training-free
@@ 4.1 Quantization-Aware Training
[…]
@@ 4.2 Post-Training Quantization
OMPQ [ma2023ompq] */44 $4.5$ $68.69$ $64$
ZeroQ [cai2020zeroq] 44/44 $5.81$ $21.20$ -
$\text{BRECQ}^{\dagger}$  [li2021brecq] 44/44 $5.81$ $69.32$ -
PACT [choi2018pact] 44/44 $5.81$ $69.20$ -
HAWQ-V3 [yao2021hawq] 44/44 $5.81$ $68.45$ -
OMPQ [ma2023ompq] */88 $4.0$ $69.41$ $64$
EMQ [dong2023emq] */8 4.0 69.92 64
@@ 4.3 Generalization of TAP to Larger Transformer
EMA [jacob2018quantization] 88/88 $75$ $30.30$ $78.82$ $28.00$
Percentile [li2019fully] 88/88 $75$ $46.69$ $78.37$ $40.93$
OMSE [choukroun2019low] 88/88 $75$ $73.39$ $79.57$ $48.55$
@@ 4.5 Ablation studies
4.6 Additional Evaluation

### v2

@@ 3.3 Evolutionary Framework of the TAP
Step 0: Initialization. At time step 0, an (initialization, empty context) prompt is input to the LLM ($\mathcal{L}$) to generate the initial proxy population $\mathcal{P}_{0}=\{f_{1},\dots,f_{N}\}$. Each proxy $f_{i}$ must include three components: 1. Natural language reasoning logic for sensitivity scoring (e.g., "convolutional channel sensitivity is proportional to per-channel FLOPs or kernel area, adjusted by feature-map resolution"); 2. Executable code for computing channel-level/layer-level sensitivity scores; 3. A bit-width allocation function that maps sensitivity scores and target compression rates to specific bit-widths (e.g., 2/4/8 bits for convolutional channels, 4/8 bits for linear layers).
Initialize template weights $w^{(0)}$ uniformly and compute $p^{(0)}$ using Eq. (8).
Step 1: Proxy Generation. At generation $t$, a template $T\sim\text{Categorical}(p^{(t)})$ and an operation type (initialization/mutation/crossover) are chosen. If $\text{op}=\text{initialization}$, we set $\mathcal{C}_{t}=\emptyset$ and invoke $\mu_{\text{init},\emptyset}$ to inject novel proxies; otherwise, we sample a context window $\mathcal{C}_{t}\subseteq\mathcal{P}_{t}$ (containing the logic and performance of historically high-fitness proxies) and invoke $\mu_{\text{op},\mathcal{C}_{t}}$. The pair $(T,\mathcal{C}_{t})$ is then fed into $\mathcal{L}$ to generate the candidate proxy set $\mathcal{P}_{t}^{\prime}$.
@@
Step 3: DPO Strategy Update. 1. Aggregate per-template fitness: Compute $\mathrm{Fitness}(T_{t})$ for each $T_{t}\in\mathcal{T}$ (e.g., average $\phi$ over candidates generated by $T_{t}$). 2. Update weights and probabilities: Apply Eq. (7) to obtain $w^{(t+1)}$ and normalize via Eq. (8) to obtain $p^{(t+1)}$. 3. Update the context window: Add the logic and performance of the top-10% high-fitness proxies in $\mathcal{P}_{t}\cup\mathcal{P}_{t}^{\prime}$ to the context library, providing better references for the next generation of proxy generation.
@@
PACT [1] 69.8 5/5 7.2 - 83.86 Fixed-precision
LSQ [15] 67.6 2/2 - - - Fixed-precision
PDQ [16] 65.0 -/- - - - Fixed-precision
Hybrid-Net* [17] 62.7 $\ast$$\ast - - -$ Post-training Deterministic
HAQ [2] 70.4 $\ast$/32 5.8 - 87.47 RL-based
HAWQ [3] 68.5 $\ast$$\ast -$ 15.6 - Sensitivity
HAWQ-V3 [4] 70.4 $\ast$$\ast$ 7.0 - 84.31 Sensitivity
DNAS [18] 70.0 -/- 6.8 - 84.75 Differentiable
EdMIPS [5] 65.9 $\ast$$\ast$ 4.7 9.5 89.46 Differentiable
EdMIPS-C [5] 59.1 $\ast$$\ast$ 4.5 0.6 89.91 Differentiable
FracBits-SAT [8] 70.6 $\ast$$\ast$ 5.81 - 86.97 Differentiable
GMPQ-C [19] 69.9 $\ast$$\ast$ 4.1 0.6 90.81 Differentiable
SMPQ [20] 72.6 $\ast$/8 - 4.9 - Differentiable
ASGA [21] 66.4 $\ast$$\ast -$ 0.6 - Differentiable
OMPQ [10] 72.08 $\ast$/5 6.7 0.45 84.98 Training-free
EMQ [22] 72.28 $\ast$/6 6.7 0.51 84.98 Training-free
TAP-C 72.63 $\ast$/8 7.5 0.42 83.18 Training-free
@@
PACT [1] 76.70 4/4 16.0 - 83.64 Fixed-precision
HAQ [2] 75.48 $\ast$/32 9.62 - 90.16 RL-based
LQ-net [23] 76.40 4/32 13.1 - 84.7 Differentiable
HAWQ-V3 [4] 75.39 $\ast$$\ast$ 18.7 - 78.1 Differentiable
@@
Onebit-width [24] 76.70 $\ast$/8 12.3 - 87.42 Differentiable
SMPQ [20] 76.2 $\ast$$\ast$ 12.4 - 87.32 Differentiable
ASGA [21] 71.5 $\ast$$\ast -$ 0.6 - Differentiable
OMPQ [10] 76.28 $\ast$/5 18.7 0.46 80.88 Training-free
EMQ [22] 76.70 $\ast$/6 17.9 0.52 81.70 Training-free
TAP-C 76.72 $\ast$$\ast$ 18.6 0.43 80.98 Training-free
@@ 4.1 Quantization-Aware Training
To validate the effectiveness of TAP, we perform quantization-aware training using TAP on ResNet-18/50 in ImageNet dataset (as depicted in Table 3). As shown in Table 3, our proposed TAP-C attains the best Top-1 accuracy on ResNet18 and slightly higher accuracy on ResNet50, while delivering the shortest search time among competing MPQ methods. On ResNet18, TAP-C reaches 72.63% Top-1 accuracy, surpassing recent training-free methods such as EMQ (72.28%) and OMPQ (72.08%), with a search cost of 0.42 GPU hours—lower than OMPQ (0.45) and EMQ (0.51). On ResNet50, TAP-C obtains 76.72% Top-1 accuracy—slightly higher than EMQ (76.70%) and higher than OMPQ (76.28%)—with the lowest search cost of 0.43 GPU hours (vs. 0.52 and 0.46, respectively). These results demonstrate that TAP-C can efficiently discover near-optimal quantization policies with minimal computational cost.
@@ 4.2 Post-Training Quantization
OMPQ [10] */44 $4.5$ $68.69$ $64$
ZeroQ [26] 44/44 $5.81$ $21.20$ -
$\text{BRECQ}^{\dagger}$  [27] 44/44 $5.81$ $69.32$ -
PACT [1] 44/44 $5.81$ $69.20$ -
HAWQ-V3 [4] 44/44 $5.81$ $68.45$ -
OMPQ [10] */88 $4.0$ $69.41$ $64$
EMQ [22] */8 4.0 69.92 64
@@ 4.3 Generalization of TAP to Larger Transformer
[…]
EMA [35] 88/88 $75$ $30.30$ $78.82$ $28.00$
Percentile [36] 88/88 $75$ $46.69$ $78.37$ $40.93$
OMSE [37] 88/88 $75$ $73.39$ $79.57$ $48.55$
@@ 4.5 Ablation studies
The impact of calibration batch randomness: To assess the robustness of the calibration step, we draw five random calibration mini-batches (16 samples each) from disjoint ImageNet categories under different seeds. As shown in Table 10, Top-1 varies only marginally across batches for both ResNet18 and MobileNetV2, indicating a stable and robust calibration process.
Table 10: Ablation: calibration batch robustness on ImageNet (Top-1 %).
Model Sample 1 Sample 2 Sample 3 Sample 4 Sample 5
ResNet18 70.21 70.17 70.14 70.26 70.24
MobileNetV2 71.71 71.81 71.75 71.77 71.68
The impact of population size $N$: We vary the evolutionary population size $N\in\{2,5,8,10\}$ and report Top-1 in Table 11. Performance remains stable across a wide range of $N$, suggesting that TAP is insensitive to this hyper-parameter.
Table 11: Ablation: sensitivity to population size NN (Top-1 %).
Model N=2 N=5 N=8 N=10
ResNet18 69.87 70.18 70.26 69.13
MobileNetV2 70.79 71.42 71.81 70.27
Code executability during evolution: We track the fraction of non-executable proxy code over generations. Table 12 shows that the proportion of inoperable candidates drops quickly as evolution proceeds, confirming that executability steadily improves while invalid programs are down-weighted by the controller.
Table 12: Ablation: non-executable code proportion by iteration range (%).
Model 1–10 11–20 21–30
ResNet18 40 20 10
MobileNetV2 30 20 10